The directions to the question: find the equation of the line parallel to the line 5=-10x-10y that passes through the point (-2,1). I need the answer to this question which i'm having trouble to find. Please help answer and thank you for your time.

the answer to tis question is Y=X+3
How to do?
from the previous eqn 5=10x-10y
Let 10y=10x-5
Becos u have to find gardient from thids eqn
Y=MX+C
10y=10x-5 (divide by 5)
2y=2x-1
bring the 2 over by dividing by 2
y=x+ 1/2
So from here u know that gradient is 1
x= -2 and y=1
by sub the 2 value
1=(-2)(1)+c
C=3
so the eqn is Y=x+3

Answer by feb feb(1)


the answer to tis question is Y=X+3
How to do?
from the previous eqn 5=10x-10y
Let 10y=10x-5
Becos u have to find gardient from thids eqn
Y=MX+C
10y=10x-5 (divide by 5)
2y=2x-1
bring the 2 over by dividing by 2
y=x+ 1/2
So from here u know that gradient is 1
x= -2 and y=1
by sub the 2 value
1=(-2)(1)+c
C=3
so the eqn is Y=x+3